# JAX-MPM: A Learning-Augmented Differentiable Meshfree Framework for GPU-Accelerated Lagrangian Simulation and Geophysical Inverse Modeling

### v1

@@ 2.2 Implementation of material point method
As shown in Fig. 1, each time step in the MPM solution procedure consists of three computational stages: particle-to-grid (P2G) transfer, grid-based updates, and grid-to-particle (G2P) transfer. These stages are described in detail below.
Figure 1: Illustration of the Material Point Method (MPM) procedure. The procedure consists of four stages: P2G transfer, grid based operations, G2P transfer, and particle state update. Black dots denote the fixed background grid, blue stars represent particles, and purple stars indicate updated particles. Dashed orange arrows illustrate interpolation contributions, while solid black arrows indicate velocity vectors.
2.2.1 Particle-to-grid transfer
In the P2G phase, particle information is projected onto the background grid using interpolation kernels. The mass at grid node $i$ is computed via a weighted sum over particles
@@ 2.2.1 Particle-to-grid transfer
The nodal force $\bm{f}_{i}^{n}$ includes internal stresses, body forces, and externally applied forces:
$\bm{f}_{i}^{n}=-\sum_{p}V_{p}^{n}\nabla\phi_{ip}^{n}\cdot\bm{\sigma}_{p}^{n}+\sum_{p}\phi_{ip}^{n}m_{p}\bm{b}_{p}+\overline{\bm{f}}_{i},$ (12)
where $V_{p}^{n}$ and $\bm{\sigma}_{p}^{n}$ are the volume and stress of particle $p$, $\bm{b}_{p}$ is the body force per unit mass, and $\overline{\bm{f}}_{i}$ represents externally applied nodal forces, defined in Eq. (8).
2.2.2 Grid velocity update
With the nodal masses and forces being obtained, the grid acceleration is calculated using Newton’s second law:
@@ 2.2.2 Grid velocity update
Boundary conditions are imposed explicitly on the grid. For instance, to enforce a fixed (Dirichlet) boundary condition on a set of nodes $\Gamma_{D}$, we impose
$\bm{a}_{i}^{n}=\bm{0},\quad\bm{v}_{i}^{n}=\bm{0},\quad\text{for }i\in\Gamma_{D}.$
2.2.3 Grid-to-particle transfer
Following the update of grid velocities, information is transferred back to the particles to advance the simulation. JAX-MPM supports several established transfer schemes, including fluid-implicit particle (FLIP) [13], particle-in-cell (PIC) [40], affine-PIC (APIC) [55] and Taylor-PIC (PIC) [71]. In this section, we describe the commonly used FLIP and PIC methods, additional variants are detailed in B.
@@ 2.2.3 Grid-to-particle transfer
PIC replaces the particle velocity with the interpolated grid velocity, introducing numerical dissipation into the system [40]. In constrst, FLIP only adds the increment of grid velocity, which preserves particle-level quantities and reduces numerical dissipation, but may introduce noise and instability [13, 28]. To balance these effects, a blended scheme [104] can be used.
@@
and the velocity gradient at each particle is evaluated by:
@@
Together, these three stages complete one full MPM time step and constitute the core computational cycle of the JAX-MPM framework.
Remark 1 (Stress Update Schemes). A critical aspect of MPM implementation is the choice of stress update scheme, which defines when stresses are evaluated relative to the solution of the momentum equation. Two common stress update schemes in MPM are Update Stress Last (USL) and Update Stress First (USF) [7]. USL updates stresses after computing internal forces and advancing velocities, offering better stability and modularity. USF updates stresses earlier, during the P2G step, which can improve energy conservation but may be less stable in dynamic problems. The implementation described above follows the USL approach. The specific stress update procedure is detailed in the next section.
@@ 2.3 Constitutive model
JAX-MPM supports constitutive models tailored for large-deformation geophysical simulations. In this study, we consider two representative materials: (1) Newtonian fluids, which are employed in dam-break scenarios, and (2) elastoplastic constitutive model with Drucker–Prager yield criterion, widely employed to capture shear and tensile failure in granular materials. The stress update procedures for each materiel models are detailed below.
@@
    where $c_{n}$ is the numerical sound speed, $\rho_{0}$ is the reference density of fluid, and $\rho_{p}^{n+1}$ is the updated fluid density at time step $(n+1)$, computed as $\rho_{p}^{n+1}=\rho_{p}^{n}/(1+\operatorname{det}(\Delta\bm{d}_{p})$.
@@
    For granular flow simulations, we adopt a non-associated Drucker–Prager (D-P) plasticity model with a tension cutoff, following [52, 73]. To maintain objectivity under large deformation, the Jaumann stress rate is used. Thus, the stress is updated by:
@@ 3 JAX-MPM: Differentiable meshfree solver for forward and inverse modeling
JAX-MPM is fully implemented within the JAX computational framework [14], a modern backend explicitly designed for differentiable programming and GPU acceleration. This design enables not only high-performance forward particle-based simulations, but also efficient gradient-based optimization for inverse modeling, data assimilation, and seamless integration with machine learning models for hybrid physics–ML workflows.
@@ 3.1 Gradient computation by automatic differentiation
To enable efficient gradient-based optimization, JAX-MPM formulates the simulation update rule as a differentiable operator, allowing gradients to propagate through the full trajectory using reverse-mode automatic differentiation.
Let a set $\mathbf{S}^{t}=\left\{\bm{x}_{p}^{t},\bm{v}_{p}^{t},\bm{F}_{p}^{t},\rho_{p}^{t}\right\}_{p=1}^{N_{p}}\in\mathcal{S}$ be a collection of structured simulation states at time step $t$, where each material point is characterized by its position $\bm{x}_{p}^{t}\in\mathbb{R}^{d}$, velocity $\bm{v}_{p}^{t}\in\mathbb{R}^{d}$, deformation gradient $\bm{F}_{p}^{t}\in\mathbb{R}^{d\times d}$, and density $\rho_{p}^{t}\in\mathbb{R}$. The space $\mathcal{S}$ denotes the appropriate vector space for these simulation states. We note that additional quantities, such as stress $\bm{\sigma}_{p}^{t}\in\mathbb{R}^{d\times d}$ and equivalent plastic strain $\varepsilon_{eq,p}^{t}\in\mathbb{R}$, may be included as needed but are omitted here for simplicity.
In the JAX-MPM solver, the evolution of the simulation states $\mathbf{S}^{t}$ is defined by a recursive mapping:
$\mathbf{S}^{t+1}=\Phi(\mathbf{S}^{t},\bm{\alpha})$ (22)
which generates the forward simulation trajectory:
$\mathbf{S}^{1}\xrightarrow{\Phi(\bm{\alpha})}\mathbf{S}^{2}\xrightarrow{\Phi(\bm{\alpha})}\cdots\xrightarrow{\Phi(\bm{\alpha})}\mathbf{S}^{t+1}$ (23)
Here, the operator $\Phi:\mathcal{S}\times\mathbb{R}^{m}\rightarrow\mathcal{S}$ denotes the one-step state update procedure of the JAX-MPM solver (see Eqs. (7)–(9)), and $\bm{\alpha}\in\mathbb{R}^{m}$ denotes the set of system parameters governing the dynamics. As described in Section 2.2, the update operator $\Phi$ is composed of three substeps: particle-to-grid (P2G) transfer $\Phi_{\text{P2G}}$, grid-based update $\Phi_{\text{U}}$, and grid-to-particle (G2P) projection $\Phi_{\text{G2P}}$, such that $\Phi=\Phi_{\text{G2P}}\circ\Phi_{\text{U}}\circ\Phi_{\text{P2G}}.$
Over a complete trajectory with $N_{t}$ time steps, the ensemble of simulation states $\mathcal{X}=\{\mathbf{S}^{t}\}_{t=1}^{N_{t}}$ satisfy the discretized governing equations (3), expressed in compact form as:
@@
Within the JAX-MPM framework, the time-stepping operator $\Phi$ is differentiable, enabling exact and efficient gradient computation with respect to the physical parameters $\bm{\alpha}$ via reverse-mode automatic differentiation. Consider an arbitrary state variable at time step $(t+1)$, denoted by $\bm{z}^{t+1}\in\mathbf{S}^{t+1}$ (e.g., selecting $\bm{z}^{t+1}:=\bm{v}^{t+1}$). Its derivative with respect to $\bm{\alpha}$ can be computed recursively through time by applying the chain rule:
@@
where $\Phi_{\bm{z}}^{t}$ denotes the time-stepping operator at step $t$, evaluated using input $\mathbf{S}^{t}$ and restricted to the output $\bm{z}$. In practice, this process continues backward through all prior steps to the initial condition via the recursive mapping in Eq. (22), e.g., $\frac{d\mathbf{S}^{t}}{d\bm{\alpha}}=\frac{\partial\Phi^{t-1}}{\partial\mathbf{S}^{t-1}}\cdot\frac{d\mathbf{S}^{t-1}}{d\bm{\alpha}}+\frac{\partial\Phi^{t-1}}{\partial\bm{\alpha}}$. This recursive structure is automatically handled by JAX’s autodiff engine, i.e., $\frac{d\bm{z}^{t+1}}{d\bm{\alpha}}=\texttt{jax.grad}(\bm{z}^{t+1})(\bm{\alpha}),$ which constructs the forward computational graph during execution and traverses it backward to compute gradients.
@@ 3.2 PDE-constrained optimization for inverse modeling
Inverse modeling plays a central role in geophysical hazard analysis, where key physical parameters—such as initial velocities or spatially varying friction—cannot be directly observed but significantly influence flow behavior and deposition in events like landslides, dam breaks, and debris flows. To recover such quantities, we formulate the task as a PDE-constrained optimization problem, where the parameters $\bm{\alpha}$ are inferred by minimizing the discrepancy between simulated and observed data, subject to physical governing equations implemented via JAX-MPM.
To this end, we define a general differentiable output mapping operator $\mathcal{Q}$ that extracts simulated quantities corresponding to observed state data $\bm{z}_{l,t}^{\text{obs}}\in\mathbb{R}^{d}$ at spatial index $l$ and time $t$. This operator enables the JAX-MPM framework to accommodate both Lagrangian (particle-based) and Eulerian (region-based) forms of supervision. The simulated counterpart at index $l$ is denoted $\mathcal{Q}_{l}(\bm{z}^{t})$, where $\bm{z}^{t}\in\mathbf{S}^{t}$ represents the model state corresponding to the observation.
For Lagrangian supervision, the index $l$ corresponds to a tracked particle, and it simply extracts the quantity associated with $l$-th particle:
@@
For Eulerian supervision, the index $l$ corresponds to a fixed spatial coordinate $\mathbf{x}_{l}$, associated with a compact surrounding region $\Omega_{l}\subset\mathbb{R}^{d}$ centered at $\mathbf{x}_{l}$. In this setting, the simulated quantity is defined as the average over all particles within this region:
$\mathcal{Q}_{l}(\bm{z}^{t})=\frac{1}{|\mathcal{P}_{l}^{t}|}\sum_{p\in\mathcal{P}_{l}^{t}}\bm{z}_{p}^{t},\quad\text{where }\mathcal{P}_{l}^{t}=\{p\mid\bm{x}_{p}^{t}\in\Omega_{l}\},$ (27)
where $\mathcal{P}_{l}^{t}$ denotes the set of material points located within the Eulerian region $\Omega_{l}$. This unified formulation provides a consistent interface for extracting outputs for different observational modalities.
Figure 2: Schematic illustration of the inverse modeling framework in JAX-MPM. A neural network $\mathcal{N}_{\theta}(\mathbf{x})$ approximates the unknown spatially varying parameter field $\alpha(\mathbf{x})$, which serves as input to the differentiable simulator. The simulation generates state trajectories $\bm{z}^{t}(\theta)$, from which observable quantities $\mathcal{Q}_{l}(\bm{z}^{t}(\theta))$ are extracted and compared against observed data $\bm{z}^{\mathrm{obs}}_{l,t}$. The loss is computed over all space-time points and backpropagated through the entire pipeline via automatic differentiation to update the neural network parameters $\theta$.
@@
Here, $\bm{z}_{l,t}^{\text{obs}}\in\mathbb{R}^{d}$ denotes the observed quantity at spatial index $l$ and time $t$, obtained from measurements or ground-truth data, while $\bm{z}^{t}$ represents the corresponding simulated quantity used for loss evaluation. $\mathcal{T}_{t}\subseteq\{1,\dots,N_{t}\}$ is the set of selected observation time steps, and $\mathcal{T}_{x}$ denotes the set of spatial indices used for supervision. Specifically, the total number of particles is denoted by $N_{p}$, and the total number of monitor regions is denoted by $N_{m}$. Then, $\mathcal{T}_{x}\subseteq\{1,\dots,N_{p}\}$ in the Lagrangian setting and $\mathcal{T}_{x}\subseteq\{1,\dots,N_{m}\}$ in the Eulerian setting. The binary mask $m_{l,t}\in\{0,1\}$ indicates the availability of ground-truth data, and the discretized physical constraints $\mathcal{J}(\mathcal{X},\bm{\alpha})=0$ are implicitly enforced by the differentiable time-stepping operator $\Phi$, as defined in (22), over the entire simulation trajectory.
In classical adjoint-based PDE-constrained optimization, gradients with respect to $\bm{\alpha}$ require deriving backward recurrence relations involving the simulation operator $\Phi$. This involves solving an adjoint system backward in time.
In our setting, the total derivative of the loss with respect to $\bm{\alpha}$ is expressed as:
$\frac{d\mathcal{L}}{d\bm{\alpha}}=\sum_{t\in\mathcal{T}_{t}}\sum_{l\in\mathcal{T}_{x}}m_{l,t}\left(\mathcal{Q}_{l}(\bm{z}^{t})-\bm{z}_{l,t}^{\text{obs}}\right)^{\top}\cdot\frac{d\mathcal{Q}_{l}(\bm{z}^{t})}{d\bm{\alpha}},$ (29)
with the sensitivity of outputs governed by the chain rule
$\frac{d\mathcal{Q}_{l}(\bm{z}^{t})}{d\bm{\alpha}}=\frac{\partial\mathcal{Q}_{l}}{\partial\bm{z}^{t}}\cdot\frac{\partial\bm{z}^{t}}{d\bm{\alpha}},$ (30)
where $\frac{d\bm{z}^{t}}{d\bm{\alpha}}$ follows the recursive structure as shown in Eq. (25).
Rather than implementing adjoint solvers manually, JAX-MPM leverages reverse-mode automatic differentiation (AD), as detailed in Section 3.1. The entire gradient is computed automatically and efficiently via
@@
This eliminates the need for Jacobian construction or manual derivation of adjoint solvers, making it particularly effective for dynamic inverse problems involving nonlinear physics and complex state trajectories.
Finally, we remark that the parameter field $\bm{\alpha}$ can take various forms depending on the specific inverse modeling problem. It may represent discrete values specified at particular time steps (e.g., global initial velocity), span the entire simulation trajectory (e.g., basal friction or material constants), or vary continuously in space and/or time (e.g., $\bm{\alpha}=\bm{\alpha}(\mathbf{x},t)$).
When the parameter field is spatially or temporally varying, directly optimizing over high-dimensional fields can be computationally challenging. To address this, we propose to implicitly represent the parameter field using a neural network $\mathcal{N}_{\theta}$, parameterized by weights $\bm{\theta}\in\mathbb{R}^{N_{\theta}}$, as illustrated in Fig. 2. Specifically, given spatial and/or temporal coordinates $\mathbf{x}$, $t$, the neural network defines the field as $\bm{\alpha}(\mathbf{x},t)=\mathcal{N}_{\theta}(\mathbf{x},t)$. This approach enables the learning of a smooth, continuous representation of the parameter field in a flexible and differentiable manner [41, 32]. In such cases, the inverse problem defined in Eq. (28) does not directly optimize the parameters $\bm{\alpha}$, but instead optimizes the neural network weights $\bm{\theta}$. Regardless of the representation, gradients with respect to either $\bm{\alpha}$ or $\bm{\theta}$ are computed efficiently and exactly through the automatic differentiation process.

### v2

@@ 2.2 Implementation of material point method
As shown in Fig. 1, each time step in the MPM solution procedure consists of three computational stages: particle-to-grid (P2G) transfer, grid-based updates, and grid-to-particle (G2P) transfer. The computation procedures at three stages are summarized below.
Particle-to-grid (P2G) transfer
In the P2G phase, particle information is projected onto the background grid using interpolation kernels. The mass at grid node $i$, consistent with the lumped mass matrix, is computed via a weighted sum over particles
@@
The nodal force vector $\bm{f}_{i}^{n}$ consists of the internal and external force vectors defined in Eq. (8):
$\bm{f}_{i}^{n}=-\sum_{p}V_{p}^{n}\nabla\phi_{ip}^{n}\cdot\bm{\sigma}_{p}^{n}+\sum_{p}\phi_{ip}^{n}m_{p}\bm{b}_{p}+\bar{\bm{f}}_{i}^{t},$ (12)
where $V_{p}^{n}$ and $\bm{\sigma}_{p}^{n}$ are the volume and stress of particle $p$, $\bm{b}_{p}$ is the body force per unit mass, and $\bar{\bm{f}}_{i}^{t}$ represents the nodal force contributed by the traction, as defined in Eq. (8). The constitutive model to compute $\bm{\sigma}_{p}^{n}$ is given in Section 2.3.
Grid-based updates
With the nodal representation of mass and force at node $i$, the grid acceleration is given as:
@@
In explicit MPM, Dirichlet conditions are typically imposed by overwriting boundary‐node velocities. However, with higher-order kernels (e.g., quadratic B-splines), particle supports are truncated at the domain edge, which breaks partition of unity and leads to spurious stress oscillations and potential instabilities [27]. To address this issue, two approaches can be used: (i) adding ghost cells so that boundary-adjacent particles retain a complete stencil [27], thereby restoring partition of unity (this approach is adopted in the present work); and (ii) applying a local kernel-correction near boundaries to enforce the partition of unity [70].
Grid-to-particle (G2P) transfer
Following the update of grid velocities, information is transferred back to the particles to advance the simulation. JAX-MPM supports several established transfer schemes, including fluid-implicit particle (FLIP) [12], particle-in-cell (PIC) [40], affine-PIC (APIC) [54] and Taylor-PIC (TPIC) [70]. In this section, we describe the commonly used FLIP and PIC methods; additional variants are detailed in B.
@@
PIC replaces the particle velocity with the interpolated grid velocity, introducing numerical dissipation into the system [40]. In contrast, FLIP only adds the increment of grid velocity, which preserves particle-level quantities and reduces numerical dissipation, but may introduce noise and instability [12, 27]. To balance these effects, a blended scheme [103] can be used.
@@
and the velocity gradient at each particle (e.g., used to update the stress tensor) is evaluated by:
@@
At the end of the G2P stage, the quantities, such as velocity, are reset before P2G stage for next time step.
Remark on stress update schemes
A critical aspect of MPM implementation is the choice of stress update scheme, which defines when stresses are evaluated relative to the solution of the momentum equation. Two common stress update schemes in MPM are update stress last (USL) and update stress first (USF) [7]. USL updates stresses after computing internal forces and advancing velocities, offering better stability and modularity. USF updates stresses earlier, during the P2G step, which can improve energy conservation but may be less stable in dynamic problems. The implementation described above follows the USL approach.
@@ 2.3 Constitutive model
JAX-MPM supports constitutive models tailored for large-deformation geophysical flows. In this study, we consider two representative materials: (1) Newtonian fluids used in the dam-break scenarios, and (2) elastoplastic solids governed by a Drucker–Prager (D–P) yield criterion with a tensile cutoff, commonly employed for granular media. The stress update procedures for each material model are detailed below.
@@
    where $c_{n}$ is the numerical sound speed, $\rho_{0}$ is the reference density of fluid, and $\rho_{p}^{n+1}$ is the updated fluid density at time step $(n+1)$, computed as $\rho_{p}^{n+1}=\rho_{p}^{n}/(1+\operatorname{det}(\Delta\bm{d}_{p}))$.
@@
    For granular flow, we adopt a non-associated Drucker–Prager (D-P) plasticity model with a tension cutoff [51, 72]. To maintain objectivity under large deformation, we employ the Jaumann objective stress rate. The stress update over a time step is written as:
@@ 3 JAX-MPM: Differentiable meshfree solver for forward and inverse modeling
JAX-MPM is implemented within the JAX framework [13], a modern system for differentiable programming with XLA-based GPU/TPU acceleration. This design enables high-performance forward particle-based simulations while providing efficient gradient-based optimization capabilities for inverse modeling and data assimilation. Moreover, the solver can integrate seamlessly with network–based machine learning (ML) models, facilitating hybrid physics–ML workflows.
@@ 3.1 Gradient computation by automatic differentiation
JAX-MPM formulates the physical simulation as a composition of differentiable operators, allowing reverse-mode automatic differentiation to propagate gradients through entire trajectories.
Let a set $\mathbf{S}^{t}=\left\{\bm{x}_{p}^{t},\bm{v}_{p}^{t},\bm{F}_{p}^{t},\rho_{p}^{t}\right\}_{p=1}^{N_{p}}\in\mathcal{S}$ denote the collection of structured simulation states at time step $t$, where each material point is characterized by its position $\bm{x}_{p}^{t}\in\mathbb{R}^{d}$, velocity $\bm{v}_{p}^{t}\in\mathbb{R}^{d}$, deformation gradient $\bm{F}_{p}^{t}\in\mathbb{R}^{d\times d}$, and density $\rho_{p}^{t}\in\mathbb{R}$. The space $\mathcal{S}$ denotes the appropriate vector space for these simulation states. We note that additional quantities, such as stress $\bm{\sigma}_{p}^{t}\in\mathbb{R}^{d\times d}$ and equivalent plastic strain $\varepsilon_{eq,p}^{t}\in\mathbb{R}$, may be included as needed but are omitted here for simplicity.
The state update from time $t$ to $t+1$ in the JAX-MPM solver (see Fig. 1 and Eqs. (10)–(18)) is defined by the mapping:
$\mathbf{S}^{t+1}=\Phi(\mathbf{S}^{t};\bm{\alpha}),\qquad\Phi:\mathcal{S}\times\mathbb{R}^{m}\to\mathcal{S},$ (22)
where $\bm{\alpha}\in\mathbb{R}^{m}$ denotes the set of model parameters governing the dynamics (e.g., material properties, friction, boundary/initial conditions). As in Section 2.2, the update operator $\Phi$ consists of three stages: particle-to-grid transfer (P2G), grid-based updates (GU), and grid-to-particle transfer (G2P), such that $\Phi=\Phi_{\text{G2P}}\circ\Phi_{\text{GU}}\circ\Phi_{\text{P2G}}.$ The forward simulation trajectory over $N_{t}$ steps is thus expressed as:
$\mathbf{S}^{0}\xrightarrow{\Phi(\bm{\alpha})}\mathbf{S}^{1}\xrightarrow{\Phi(\bm{\alpha})}\cdots\xrightarrow{\Phi(\bm{\alpha})}\mathbf{S}^{N_{t}}$ (23)
Over a complete trajectory with $N_{t}$ time steps, the ensemble of simulation states is given by $\mathcal{X}=\{\mathbf{S}^{t}\}_{t=0}^{N_{t}}$. Accordingly, the discrete dynamics equations (23) on $\mathcal{X}$ can be expressed compactly as:
@@
where $\mathcal{J}$ denotes the operator that enforces the physical constraints.
By making the time-stepping operator $\Phi$ differentiable within the JAX-MPM framework, efficient gradient computation with respect to the physical parameters $\bm{\alpha}$ can be obtained via reverse-mode automatic differentiation. Consider a state variable of interest extracted from the full state set $\mathbf{S}$ at time step $(t+1)$, $\bm{z}^{t+1}\in\mathbf{S}^{t+1}$, for example, velocity $\bm{z}^{t+1}:=\bm{v}^{t+1}$. Its sensitivity with respect to the parameters $\bm{\alpha}$ can be evaluated recursively through time using the chain rule:
@@
where the subscript $z$ in $\Phi_{\bm{z}}^{t}$ indicates that the component $\bm{z}$ is extracted from the output of the time-stepping operator evaluated using input $\mathbf{S}^{t}$. The recursion in (25) continues backward in time via Eq. (22), e.g.,
$\frac{d\mathbf{S}^{t}}{d\bm{\alpha}}=\frac{\partial\Phi^{t-1}}{\partial\mathbf{S}^{t-1}}\cdot\frac{d\mathbf{S}^{t-1}}{d\bm{\alpha}}+\frac{\partial\Phi^{t-1}}{\partial\bm{\alpha}}.$ (26)
In practice, this recursion is handled automatically by JAX’s autodiff engine. For example, consider a scalar quantity $z^{t+1}$ that depends on $\bm{\alpha}$. Its gradient can be obtained by invoking jax.grad
$\frac{dz^{t+1}}{d\bm{\alpha}}=\texttt{jax.grad}(z^{t+1})(\bm{\alpha}).$ (27)
JAX constructs the computation graph during the forward pass and subsequently evaluates gradients by traversing it backward using reverse-mode vector–Jacobian products, thereby eliminating the need for explicit recursive implementation.
@@ 3.2 PDE-constrained optimization for inverse modeling
Inverse modeling plays a central role in geophysical hazard analysis, where key physical parameters, such as initial velocities or spatially varying friction, cannot be directly observed but significantly influence flow behavior and deposition in events like landslides, dam breaks, and debris flows. We formulate parameter estimation as a PDE-constrained optimization problem, in which the unknown parameters $\bm{\alpha}$ are inferred by minimizing the discrepancy between simulated and observed data, subject to physical governing equations implemented within JAX-MPM.
To this end, we introduce a differentiable output mapping operator $\mathcal{Q}$ that extracts simulated quantities corresponding to observed state data $\bm{z}_{l,t}^{\text{obs}}\in\mathbb{R}^{d}$ at spatial index $l$ and time $t$. This operator provides a unified interface that accommodates both Lagrangian (particle-based) and Eulerian (region-based) forms of supervision. The simulated counterpart at index $l$ is denoted $\mathcal{Q}_{l}(\bm{z}^{t})$, where $\bm{z}^{t}\in\mathbf{S}^{t}$ represents the model state associated with the observation.
  • 1.
    In the Lagrangian setting, $l$ denotes a tracked particle, and the operator simply returns its state:
@@
  • 2.
    In the Eulerian setting, $l$ corresponds to a fixed spatial coordinate $\mathbf{x}_{l}$, associated with a compact surrounding region $\Omega_{l}\subset\mathbb{R}^{d}$ centered at $\mathbf{x}_{l}$. The simulated output is defined as the average over particle states within this region:
    $\mathcal{Q}_{l}(\bm{z}^{t})=\frac{1}{|\mathcal{P}_{l}^{t}|}\sum_{p\in\mathcal{P}_{l}^{t}}\bm{z}_{p}^{t},$ (29)
    where $\mathcal{P}_{l}^{t}=\{p\mid\bm{x}_{p}^{t}\in\Omega_{l}\}$ denotes the set of material points located within the Eulerian region $\Omega_{l}$.
Figure 2: Schematic illustration of the inverse modeling framework in JAX-MPM. A neural network $\mathcal{N}_{\theta}(\mathbf{x})$ approximates the unknown spatially varying parameter field $\alpha(\mathbf{x})$, which serves as input to the differentiable simulator. The simulation generates state trajectories $\bm{z}^{t}(\theta)$, from which observable quantities $\mathcal{Q}_{l}(\bm{z}^{t}(\theta))$ are extracted and compared against observed data $\bm{z}^{\mathrm{obs}}_{l,t}$. The loss is computed over selected space-time points and backpropagated through the entire pipeline via automatic differentiation to update the neural network parameters $\theta$.
@@
where the discrete physics $\mathcal{J}(\mathcal{X},\bm{\alpha})=\bm{0}$ are enforced implicitly by the differentiable time-stepping operator $\Phi$ in Eq. (22) along the trajectory $\mathcal{X}=\{\mathbf{S}^{t}\}_{t=0}^{N_{t}}$. The binary mask $m_{l,t}\in\{0,1\}$ indicates the availability of ground-truth data at spatial index $l$ and time $t$. $\bm{z}_{l,t}^{\text{obs}}\in\mathbb{R}^{d}$ denotes the observed quantity at spatial index $l$ and time $t$. The set of observed time indices is $\mathcal{T}_{t}\subseteq\{0,\dots,N_{t}\}$, where $N_{t}$ is the total number of time steps. The set of spatial indices is $\mathcal{T}_{x}$, which depends on the supervision type:
[…]
with $N_{p}$ denoting the number of particles and $N_{m}$ the number of monitor regions.
In classical adjoint-based PDE optimization, gradients w.r.t. $\bm{\alpha}$ require deriving and solving backward recurrences. In our setting, the loss gradient decomposes as
$\frac{d\mathcal{L}}{d\bm{\alpha}}=\sum_{t\in\mathcal{T}_{t}}\sum_{l\in\mathcal{T}_{x}}m_{l,t}\,\bigl(\mathcal{Q}_{l}(\bm{z}^{t})-\bm{z}_{l,t}^{\text{obs}}\bigr)^{\top}\frac{d\mathcal{Q}_{l}(\bm{z}^{t})}{d\bm{\alpha}},\quad\frac{d\mathcal{Q}_{l}(\bm{z}^{t})}{d\bm{\alpha}}=\frac{\partial\mathcal{Q}_{l}}{\partial\bm{z}^{t}}\,\frac{d\bm{z}^{t}}{d\bm{\alpha}},$ (32)
with $d\bm{z}^{t}/d\bm{\alpha}$ obtained recursively via Eq. (25). Rather than implementing adjoint solvers manually, JAX-MPM leverages reverse-mode automatic differentiation (AD), as detailed in Section 3.1:
@@
We remark that the parameter field $\bm{\alpha}$ can take several forms, depending on the inverse problem. It may represent (i) discrete values specified at particular time steps (e.g., an initial velocity), (ii) trajectory-wide constants (e.g., basal friction or material parameters), or (iii) a field varying in space and/or time, i.e., $\bm{\alpha}=\bm{\alpha}(\mathbf{x},t)$.
When $\bm{\alpha}$ varies spatially or temporally, direct optimization over high-dimensional fields is costly. We therefore parametrize the field with a light neural network $\mathcal{N}_{\theta}$ with weights $\bm{\theta}\in\mathbb{R}^{N_{\theta}}$ (see Fig. 2), such that
$\bm{\alpha}(\mathbf{x},t)\approx\bm{\alpha}(\mathbf{x},t;\bm{\theta})\;:=\;\mathcal{N}_{\theta}(\mathbf{x},t)$ (34)
This approach provides a smooth and generalized representation of the parameter field for efficient learning [41, 31]. In this setting, the inverse problem in Eq. (30) optimizes $\bm{\theta}$ rather than $\bm{\alpha}$; equivalently, one solves $\min_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ with the constraint $\mathcal{J}(\mathcal{X},\bm{\alpha}(\bm{\theta}))=\bm{0}$. Gradients with respect to either $\bm{\alpha}$ or $\bm{\theta}$ are computed via reverse-mode automatic differentiation through the simulator and the network.
Numerical studies demonstrating inverse modeling using both Lagrangian and Eulerian observations, as well as neural network parameterizations, are presented in Section 5.